Sigt hoy each,th residrts ola Plays spend allof t one re tings every year:they colectively buy 250 bathing auts6 tubes of sunscreen, and 400 beach towels. ltem (Amount purchased Bathing suits 250) Price Price Price Price 2006 ($) $10.00 4.00 5.00 2007 ($) $12.00 5.00 5.50 2008 (3) $15.00 5.00 7.00 2009 ($) $18.00 6.00 9.00 600) Beach towels (400) Use the data in the table above to calculate the following a. The total cost of this basket each year from 2006 through 2009 Instructions: Round your answers to the nearest whole number Price 2007 (S Price 2008 (5 2009 (5) 2006 ($) Total cost of basket b. How much has the cost of this basket changed from year to year in percentage terms? Instructions: Round your answers to one decimal place. 2006-2007 2007-2008 Percentage change in total cost of basket
Answer #1

A.

Cost of basket in 2006 = 250*10 + 600*4 + 400*6 = $7300

Cost of basket in 2007 = 250*12 + 600*5 + 400*5.5 = $8200

Cost of basket in 2008 = 250*15 + 600*5 + 400*7 = $9550

Cost of basket in 2009 = 250*18 + 600*6 + 400*9 = $11700

B.

% change in cost of basket between 2006-2007 = (8200-7300)/7300 = 12.3%

% change in cost of basket between 2007-2008 = (9550-8200)/8200 = 16.5%

% change in cost of basket between 2008-2009 = (11700-9550)/9550 = 22.5%